Problem 3-33 Adam Granger operates a kiosk in downtown Chicago, at which he sells one style of baseball hat. He buys the hats from a supplier for $30 and sells them for $36. Adams current breakeven point is 25,500 hats per year. 回Your answer is correct. Calculate contribution margin per unit. Contribution margin per unit SHOW SOLUTION SHOW ANSWER LINK TO TEXI LINK TO TEXT [ Your answer is correct. What is Adams current level of fixed costs? (Use the rounded contribution margin per unit calculated in the previous part.) 153,00 Current level of fixed costs $
media%2F214%2F214ef051-4e51-437c-8498-cf
media%2F322%2F3228c75f-f59f-4e40-9c53-98
Answer #1

Answer

1.

Lets units sold = x

Contribution Margin = Sales – Variable cost

= ($36 * x) – ($30 * x)

Contribution Margin = 6x

Net income before tax = Contribution margin – Fixed Cost

Net income before tax = 6x – 153,000

Net Income after Tax = Net income before tax – 30%

= 6x – 153,000 – [30% * (6x – 153,000)]

= 6x – 153,000 – 1.8x + 45900

Net Income after Tax = 4.2x – 107,100

Profit = Net income after tax

35,700 = 4.2x – 107,100

4.2x = 142,800

X = 34,000 Hats

Adam must have sold 34,000 Hats last year.

2.

Margin of Safety = Actual Sales – BEP (In value)

BEP (In Value) = BEP (In Units) * Sales price per Hat

= 25,500 Hats * $36 per unit

BEP (In Value) = $918,000

Actual Sales = 34,000 * $36 per Hat

Actual Sales = $1,224,000

Margin of Safety = Actual Sales – BEP (In value)

= 1,224,000 - 918,000

Margin of Safety = $306,000

3.

Lets units to be sold = x

Contribution Margin = Sales – Variable cost

= ($36 * x) – ($30 * x)

Contribution Margin = 6x

Net income before tax = Contribution margin – Fixed Cost

Net income before tax = 6x – 153,000

Net Income after Tax = Net income before tax – 30%

= 6x – 153,000 – [30% * (6x – 153,000)]

= 6x – 153,000 – 1.8x + 45900

Net Income after Tax = 4.2x – 107,100

Profit = Net income after tax

64,260= 4.2x – 107,100

4.2x = 171,360

X = 40,800 Hats

Adam have to sell 40,800 Hats to earn $64,260 Profit.

4.

New Contribution margin per hat = $5 per hat ($36 - $31)

Fixed Cost = 153,000 (Will remain same)

New BEP = Fixed Cost / new Contribution per hat

= 153,000 / 5

New BEP = 30,600 Hats

5.

New Sale price

37

Variable Cost

31

Units (34,000 – 5%)

32300

Sales

1,195,100

Variable Cost

1,001,300

Contribution Margin

193,800

Fixed Cost

153,000

Net profit before tax

40,800

Tax @ 30%

12,240

Net Profit after tax

28,560